OpenStudy (anonymous):

i have a cylinder with a volume of 603.1857895, and surface area is 351.8583772, i need to find the radius and the height

OpenStudy (anonymous):

Volume of a cylinder = (pi) * r^2 = 603.1857895 Surface Area of a Cylinder = 2(pi) * r^2 + 2(pi) * r * h = 351.8583772 is that good?

OpenStudy (anonymous):

ops, typo. Volume of a cylinder = (pi) * r^2 * h = 603.1857895 Surface Area of a Cylinder = 2(pi) * r^2 + 2(pi) * r * h = 351.8583772 two unknowns and two equations

OpenStudy (anonymous):

no, i already new that don't know where to go from there

OpenStudy (anonymous):

r(1)=0.103 r(2(=56,3

OpenStudy (anonymous):

i am trying to build a cylinder and all i have is the surface area and the volume, nothing else, i have been working on this for hours and very frustrated. i have found many places that show me the formulas but none that actually show you how to do it backwards

OpenStudy (anonymous):

\[Volume = (\pi)r^2h = 603.1857895\]\[Surface = 2(\pi)r^2 + 2(\pi)rh = 351.8583772 \] from volume equation:\[h = {603.1857895 \over \pi r^2} \] substitute in the surface eqt.:\[2 \pi r^2 + 2 \pi r ({603.1857895 \over \pi r^2}) = 351.8583772\]\[2 \pi r^2 + {2*603.1857895 \over r} = 351.8583772\] solve for r and then solve for h

OpenStudy (anonymous):

does that hlep?

OpenStudy (anonymous):

the actual paper i have has it written as 192 times pi, for the volume, and 112 times pi for the surface area, my teacher sucks and doesn't use 3.14 for pi, would this make this equation any easier

OpenStudy (anonymous):

it is a lot easier if yu just use the symbol (pi) instead of the actual number because at the end it is most likely to be canceled..

OpenStudy (anonymous):

for example, when i found H. If you had just left as the (pi) symbol:\[h={603.1857895 \over \pi r^2} = { 192 \pi \over \pi r^2} = { 192 \over r^2}\] See ?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

now it just made everything much easier. Can yu solve from now on?

OpenStudy (anonymous):

i think so